2x+4y=-3
2x-y=7
y=2x-7
2x+4y=-3
2x+4(2x-7)=-3
2x+8x-28=-3 can you show me step by step please? their answer was 5/2,-2?

Answer by richwmiller(17219)

So far so good.
Keep going.
2x+8x-28=-3
combine
10x=25
x=25/10
x=5/2
plug 5/2 for x in either equation
2*(5/2)-y=7
5-y=7
subtract 7 from both sides
-2-y=0
add y to both sides
-2=y